A 500 kg car drive 1 km along a horizontal road at 25 km/hrs then how much energy is required to the car move in 1 km

Grade:12th pass

1 Answers

Arun
6 years ago
 =2as => a=200 m/s2Force applied by hammer= 200x1=200 Ntime of impact= 2/200=0.01 seconds=10 milli seconds2-u2v=0 m/s , u=2 m/s => v
hope it helps